OpenStudy (anonymous):

Traveling at 55 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 310 miles from Boston to Syracuse? 6 338 365 3,300

OpenStudy (aveline):

What's 310/55? Multiply by 60 to convert into minutes

OpenStudy (texaschic101):

55/60 = 310/x....55 miles to 60 minutes = 310 miles to x minutes now cross multiply, and solve for x

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

this is confussing me really badly and I only have 5 more mins

OpenStudy (aveline):

310/55 let's you know how many hours it will take. Simply multiply by 60 to convert the hours into minutes

OpenStudy (aveline):

310/55=5.63636363 5.63636363*60=338 minutes

OpenStudy (texaschic101):

that is correct as well